OpenStudy (anonymous):

How do i use Cauchy's Integral formula with 1/(exp(z)-1) as the f(z)

OpenStudy (anonymous):

\[\int\limits_{C}^{} f(z) dz where C is \[\left| z \right| =1\] and \[f(z) = 1/(e ^{z}-1)\]